need help

Suppose the car doesn't rebound off the wall, but
Answer #1

impulse j = Favg*dt = m*(v-u)


here v is the final speed = 0 m/s

u is the initial speed = not given in the problem

m is the mass of the car

dt = 0.15 s

then average force is Favg = m*u/dt = m*u/0.15 = 6.67*m*u


if we know m and u ,we can substitute to get Favg